# Priors in Time: Missing Inductive Biases for Language Model Interpretability

### v1

@@ 5 Temporal Feature Analysis: Towards Predictive Approaches for Interpreting Neural Representations
Before analyzing how different approaches represent the temporal structure of language, we demonstrate that Temporal Feature Analysis performs on par with standard SAEs on standard metrics such as reconstruction error. Specifically, we train a Temporal Feature Analyzer and standard SAEs (ReLU, TopK, BatchTopK) on 11B token activations extracted from Gemma-2-2B (Team et al., 2024) from the Pile-Uncopyrighted dataset (Monology, 2021). We also analyze a baseline of the prediction only module from Temporal Feature Analysis, reported as ‘Pred. only’, which can be expected to underperform since predicting the next-token representation is likely to be more difficult than reconstructing it. Results are provided in Tab. 12 and show competitive performance between all protocols, except Pred. only. One can also assess which part of a fully trained Temporal Feature Analyzer is more salient in defining its performance, i.e., does the estimated predictive part $\hat{\bm{x}}_{p,t}$ contribute more to the reconstruction $\hat{\bm{x}}$ or does the estimated novel part $\hat{\bm{x}}_{n,t}$. Results are reported in Tab. 3. We see that the error vectors, i.e., $\bm{x}-\hat{\bm{x}}_{p}$ and $\bm{x}-\hat{\bm{x}}_{n}$, are approximately orthogonal, suggesting the modules computing predictive and novel codes capture separate bits of information from the input signal. Furthermore, we find that a bulk of the reconstructed signal $\hat{\bm{x}}_{t}$ (in the sense of norm) is captured by the predictive code—in fact, the percentage contribution of the predictive code is $\sim$80%, in line with numbers observed in Fig. 2d. However, analyzing the reconstruction performance, we see the predictive component primarily contributes to achieving a good NMSE, while the novel component is more responsible for explaining the input signal variance. These results align with the generative model assumed in Temporal Feature Analysis (Eq. 3). Specifically, NMSE captures the average reconstruction, and hence a slower moving, contextual signal can expect to dominate its calculation; in fact, we see the attention layer used for defining the prediction module produces block structured attention maps related to sub-event chunks (see Fig. 12), suggesting more granular temporal dynamics are being captured by predictive part (we elaborate on this last point in the next section). Meanwhile, variance assesses changes per dimension and timestamp in the signal, which better matches the inductive bias imposed on the novel part.
@@ 6 Capturing dynamic structure with Temporal Feature Analysis
We analyze standard SAEs (ReLU, TopK, and BatchTopK) and Temporal Feature Analyzers trained on 1B model activations from Layer 15 of a Gemma-2-2B model extracted from the Pile Uncopyrighted dataset (Monology, 2021). For evaluation domains, we again seek inspiration from cognitive and computational neuroscience literature on assessing humans’ abilities to process temporal structure in language. These settings capture a spectrum of interaction between a system’s prior knowledge versus temporally offered information about a concept, i.e., how the concept gets used within the given context. Specifically, we first use stories as a testbed for evaluation. Motivated by prior work on understanding how narrative structures are parsed by humans (Rumelhart, 1975; Thorndyke, 1977; Baldassano et al., 2017; Nastase et al., 2021), we first analyze whether local event structures from stories are reflected in latent codes derived out of SAEs and Temporal Feature Analysis. This help us evaluate how different protocols represent local vs. global (over time) semantic information. To investigate a more syntactic variant of this experiment, we analyze latent codes extracted from garden path sentences, i.e., grammatically valid but ambiguous phrases in which the constituents, when parsed according to their typical syntactic role, lead to an incorrect overall sentence parse (e.g., “The old man the road”). Thus, this setting captures a domain wherein the context requires interpreting constituent words in a manner that goes against the most likely interpretation from a model’s pretraining prior, i.e., the temporal information dominates the interpretation of a constituent. While humans can resolve these sentences by backtracking (Hale, 2001; Hanna and Mueller, 2024; Levy, 2008), LMs operate in an autoregressive manner and hence must capture the correct sentence parse in the forward direction. Given that LMs possess the ability to do so (Li et al., 2024; Hanna and Mueller, 2024), we compare whether latent codes extracted using SAEs and Temporal Feature Analysis relate sentence constituents according to the valid parse. Finally, motivated by literature on cognitive maps (Behrens et al., 2018), wherein a subject has to infer a latent in-context structure that relates two observations (tokens)—a capability LMs have been show to possess (Park et al., 2025)—we assess whether different interpretability protocols’ latent codes capture these in-context representations. In this last setting, a model cannot rely on its prior knowledge at all, since the behavior is entirely guided by the in-context, temporal structure.
@@ 6.1 The Geometry of Stories: A Narrative-Driven Domain
We consider the TinyStories datasets (Eldan and Li, 2023) for its relatively straightforward narrative structures, and qualitatively analyze the geometry of latent codes extracted from model activations when processing these stories. Visualizing the latent codes in a low-dimensional basis via a 3D UMAP projection (McInnes et al., 2018), we see SAEs yield a highly irregular and unstructured geometry (see Fig. 5a). Calculating Tortuosity (Bullitt et al., 2003), a measure of how aligned local arcs are with respect to the global structure of a curve, we see very high values emerge for SAEs’ latent codes geometry, suggesting sudden changes in the local similarity as a story unravels. To further understandthe results above, we highlight a specific token (‘and’) from the story, the UMAP analysis shows that standard SAEs generally just cluster tokens by lexical identity. This is further corroborated by running a hierarchical clustering algorithm on the latent codes (SciPy, 2025), finding temporally incoherent, but lexically related clusters (see Fig. 5b).
Performing the experiments above on latent codes extracted from Temporal Feature Analyzers, we see that while the novel component exhibits behavior that is similar to standard SAEs’, the predictive component shows a smooth, regular curve that clusters together several tokens from the story around a similar point in space. Again running hierarchical clustering, the extracted dendrograms suggest these clusters qualitatively correspond to events occurring in the story. We emphasize this finding is very similar to the phenomenon of temporal straightening observed in neural recordings of human subjects when processing stories (Xu et al., 2025)! Related findings show that such straightening is in fact exhibited by language models as well (Hosseini and Fedorenko, 2023), which our proposed protocol recovers in an unsupervised manner.
@@
Quantifying Similarity to Slow vs. Fast Moving Signals. To further quantify the straightening claim, we compute the Fourier transform of the model activations and divide the frequency spectrum into two halves at a critical frequency $f_{c}$ such that the energy (i.e., sum of squared value of phase information) in the frequencies below $f_{c}$ equals that of the remaining ones. We call the first split “slow part” of a sequence, and latter the “fast part”. We then compute the correlation matrix defined by the slow and fast parts, compute their spectrum, and analyze how similar these spectrums are to the ones defined using different SAEs’ codes. Results are shown in Fig. 6. We clearly see the spectrum extracted from the predictive component of Temporal Feature Analysis approximates that of the slow part of the representation, while the novel component’s spectrum is similar to that of the fast moving part. Meanwhile, spectra of SAEs only exhibits similarity to the fast part, suggesting they do not capture longer range dependencies necessary for interpreting narrative-driven texts like most language domains. To quantify this further, we also measure kernel similarities (CKA) between the kernels respectively defined by the slow moving and fast moving signal to the different SAE codes. Results clearly show Standard SAEs are only similar to fast moving part of the representation.
@@
The results above demonstrate Temporal Feature Analyzers’ predictive component qualitatively align with event boundaries in a story. To investigate this result more quantitatively, we use GPT-5 to create a synthetic dataset of 50 stories with well-defined event boundaries (see Fig. 7a for an example). We extract latent codes for these stories’ tokens, center them by subtracting the mean to remove any globally shared information, and compute the cosine similarity of token to token latent codes. If the latent codes reflect local event structure of a story, the cosine similarity (on average) will be high between token pairs that come from the same event and low (if not zero) between pairs sampled across events; see Fig. 7b for an example similarity map corresponding to the story shown in Fig. 7a. Results are shown in Fig. 7 (c) and corroborate our qualitative findings: we see predictive components from Temporal Feature Analyzers show substantially higher similarity of codes if tokens are sampled from within an event, while the novel component and SAEs generally show low similarity between two tokens. These results are further supported by the robustness of Temporal Feature Analysis to noise. Specifically, we see that when we add noise to the input data, which, on average, will lead to turning off of latents with small magnitudes (due to the encoding nonlinearity) and correspondingly amplify the temporal structure of the data, if it is present. We see precisely this effect in Fig. 7d: the predictive components’ cosine similarity map under Gaussian noised input maps elicits coarser block structures with increasing noise scale; this is reminiscent of percolation or heat diffusion perspectives on graph clustering, wherein noise diffuses only within a connected component and hence community structure is elicited by noise node features (Von Luxburg, 2007). Correspondingly, we see Temporal Feature Analyzers respond most gracefully to noise: variance explained reduces slower than SAEs’, which in fact drops to $\sim$0 at some critical noise scale.
@@ 6.2 Garden Path Sentences: Ambiguities Resolved via Temporal Structure
Garden-path sentences—e.g., The old man the boat—initially cue an incorrect local parse before a later token forces reanalysis. Language models have been shown to be able to correctly parse such ambiguous sentences, offering in fact a predictive account of human per-token surprisals (Li et al., 2024; Hanna and Mueller, 2024; Oh and Schuler, 2023). Interestingly, when using SAE codes to assess whether LLM representations offer a valid parse of the sentence, we find hierarchical clustering of SAE codes yields a parse that is suggestive of the misleading cue; meanwhile, Temporal Feature Analysis recovers the correct parse by separating the predictable, slow-moving component of the representation from the novel, fast-changing residual. Specifically, in Fig. 8, we see the predictive codes link long-distance heads and dependents that reflect the correct parse (e.g., man as verb), producing coherent similarity structure over the full span, whereas standard SAEs emphasize only transient, local changes and miss these cross-temporal constraints (see App. D.4 for more examples).
@@
Figure 9: In-context representations defined by temporal relations. (a) Schematic of the synthetic “cognitive map” task defined by Park et al. (2025): a random walk over a small graph is concatenated into a string of tokens, which when fed into a Gemma-2-2B model leads to model next-token predictions that align with the graph structure. This behavior is exhibited in the model’s activations, since a mere 2D PCA shows the graph structure. (b) Using latent codes for low-dimension projection, we see only the predictive component of the Temporal SAE clearly recovers the graph’s geometry, whereas standard SAEs and the novel component only partially reflect it. (c) Counterfactual editing: projecting the final-token representation onto the PCA basis and replacing it with a target node’s code reliably shifts next-token predictions in a graph-consistent manner.
@@ 6.3 In-Context Representations defined by entirely temporal relations
We finally use the recently described phenomenon of in-context representations by Park et al. (2025) as a domain to compare SAEs and Temporal Feature Analyzers. In particular, inspired by experiments on cognitive maps (Behrens et al., 2018), Park et al. (2025) propose to take a structured graph whose nodes are assigned entities a pretrained LM is likely to have seen before during training (e.g., apple, car, etc.); see Fig. 9a for an example. A random walk of length 500 tokens is performed on this graph and the sampled nodes during the walk are concatenated together into a string. When this is inputted into a pretrained language model, provided enough context, one finds (see Fig. 9a): (i) the model begins to produce next-token predictions inline with the graph structure and (ii) a 2D PCA projection of the model activations reflects this structure. That is, the graph structure, which is randomly defined and hence impossible for the model to have seen during pretraining, is an entirely temporal construct that is inferred by the model in-context.
@@ 7.1 Geometry of latent codes in an OOD dialogue domain
We analyze a multi-turn chat domain to test whether temporal feature analysis continues to reveal slow-moving structure. We emphasize the activations in this experiment are sampled from a Gemma-2-2B-IT model, although all analyzed interpretability protocols were trained on the base (non-instruction tuned) model. Thus, in a sense, the experiments in this section also gauge the ability of different protocols to generalize out-of-distribution (OOD)—a property SAEs have been argued to partially possess when transferring between base and the corresponding instruction-tuned models (Kissane et al., 2024). Specifically, we use the Ultrachat dataset, where we take the first 350 tokens from 1,000 conversation (mostly single-turn within this window, with up to four turns). For each token we compute a shared 2D UMAP embedding from (i) raw model activations and (ii) latent codes from ReLU, TopK, and BatchTopK SAEs, as well as the novel and predictive components of the Temporal SAE. To visualize population structure clearly, we display only 100 conversations at a time to avoid clutter (Fig. 10), though fitting is applied to all 1,000.
At the population level, the predictive component of Temporal Feature Analysis is the only representation that cleanly organizes the data along two interpretable axes: (1) a speaker-role separation, with user and assistant tokens forming distinct yet smoothly adjoining manifolds, and (2) a temporal gradient, with positions flowing continuously from early to late tokens along the manifold. By contrast, standard SAEs and the novel code show diffuse clouds with weaker role separability and no coherent positional gradient (Fig. 10, top row). To assess within-sequence geometry, we overlay a single conversation’s trajectory onto the population embedding (Fig. 10, bottom row). Predictive codes trace a smooth trajectory that remains locally coherent within each role segment and transitions gracefully at speaker switches. In contrast, standard SAEs and the novel code produce jagged, zig–zagging paths. We again quantify this with tortuosity $\tau$ (Bullitt et al., 2003) (lower is straighter): predictive codes achieve $\tau\approx 2.0$, whereas SAEs and the novel code yield $\tau\approx 25$3030, indicating substantially more local turning (Fig. 10, panel headers). This straightening mirrors the story-domain results (Sec. 6.1), now in an OOD conversational chat regime.

### v2

@@ 5 Temporal Feature Analysis: Towards Predictive Approaches for Interpreting Neural Representations
Before analyzing how different approaches represent the temporal structure of language, we demonstrate that Temporal Feature Analysis performs on par with SAEs on standard metrics such as reconstruction error. Specifically, we train a Temporal Feature Analyzer and standard SAEs (ReLU, TopK, BatchTopK) on 11B token activations extracted from Gemma-2-2B (team2024gemma) from the Pile-Uncopyrighted dataset (monology2021pile-uncopyrighted). We also analyze a baseline of the prediction only module from Temporal Feature Analysis, reported as ‘Pred. only’, which can be expected to underperform since predicting the next-token representation is likely to be more difficult than reconstructing it. Results are provided in Tab. 12 and show competitive performance between all protocols, except Pred. only. One can also assess which part of a fully trained Temporal Feature Analyzer is more salient in defining its performance, i.e., does the estimated predictive part $\hat{\bm{x}}_{p,t}$ contribute more to the reconstruction $\hat{\bm{x}}$ or does the estimated novel part $\hat{\bm{x}}_{n,t}$. Results are reported in Tab. 3. We see that the error vectors, i.e., $\bm{x}-\hat{\bm{x}}_{p}$ and $\bm{x}-\hat{\bm{x}}_{n}$, are approximately orthogonal, suggesting the modules computing predictive and novel codes capture separate bits of information from the input signal. Furthermore, we find that a bulk of the reconstructed signal $\hat{\bm{x}}_{t}$ (in the sense of norm) is captured by the predictive code—in fact, the percentage contribution of the predictive code is $\sim$80%, in line with numbers observed in Fig. 2d. However, analyzing the reconstruction performance, we see the predictive component primarily contributes to achieving a good NMSE, while the novel component is more responsible for explaining the input signal variance. These results align with the generative model assumed in Temporal Feature Analysis (Eq. 3). Specifically, NMSE captures the average reconstruction, and hence a slower moving, contextual signal can expect to dominate its calculation; in fact, we see the attention layer used for defining the prediction module produces block structured attention maps related to sub-event chunks (see Fig. 13), suggesting more granular temporal dynamics are being captured by predictive part (we elaborate on this last point in the next section). Meanwhile, variance assesses changes per dimension and timestamp in the signal, which better matches the inductive bias imposed on the novel part.
@@ 6 Capturing dynamic structure with Temporal Feature Analysis
We analyze standard SAEs (ReLU, TopK, and BatchTopK) and Temporal Feature Analyzers trained on 1B model activations from Layer 15 of a Gemma-2-2B model extracted from the Pile Uncopyrighted dataset (monology2021pile-uncopyrighted). For evaluation domains, we again seek inspiration from cognitive and computational neuroscience literature on assessing humans’ abilities to process temporal structure in language. These settings capture a spectrum of interaction between a system’s prior knowledge versus temporally offered information about a concept, i.e., how the concept gets used within the given context. Specifically, we first use stories as a testbed for evaluation. Motivated by prior work on understanding how narrative structures are parsed by humans (rumelhart1975notes; thorndyke1977cognitive; baldassano2017discovering; nastase2021narratives), we first analyze whether local event structures from stories are reflected in latent codes derived out of SAEs and Temporal Feature Analysis. This help us evaluate how different protocols represent local vs. global (over time) semantic information. To investigate a more syntactic variant of this experiment, we analyze latent codes extracted from garden path sentences, i.e., grammatically valid but ambiguous phrases in which the constituents, when parsed according to their typical syntactic role, lead to an incorrect overall sentence parse (e.g., “The old man the road”). Thus, this setting captures a domain wherein the context requires interpreting constituent words in a manner that goes against the most likely interpretation from a model’s pretraining prior, i.e., the temporal information dominates the interpretation of a constituent. While humans can resolve these sentences by backtracking (hale2001probabilistic; hanna2024incremental; levy2008expectation), LMs operate in an autoregressive manner and hence must capture the correct sentence parse in the forward direction. Given that LMs possess the ability to do so (li2024incremental; hanna2024incremental), we compare whether latent codes extracted using SAEs and Temporal Feature Analysis relate sentence constituents according to the valid parse. Finally, motivated by literature on cognitive maps (behrens2018cognitive), wherein a subject has to infer a latent in-context structure that relates two observations (tokens)—a capability LMs have been show to possess (park2025iclrincontextlearningrepresentations)—we assess whether different interpretability protocols’ latent codes capture these in-context representations. In this last setting, a model cannot rely on its prior knowledge at all, since the behavior is entirely guided by the in-context, temporal structure.
@@ 6.1 The Geometry of Stories: A Narrative-Driven Domain
We consider the TinyStories datasets (eldan2023tinystories) for its relatively straightforward narrative structures, and qualitatively analyze the geometry of latent codes extracted from model activations when processing these stories. Visualizing the latent codes in a low-dimensional basis via a 3D UMAP projection (mcinnes2018umap), we see SAEs yield a highly irregular and unstructured geometry (see Fig. 5a). Calculating Tortuosity (bullitt2003measuring), a measure of how aligned local arcs are with respect to the global structure of a curve, we see very high values emerge for SAEs’ latent codes geometry, suggesting sudden changes in the local similarity as a story unravels. To further understand the results above, we highlight a specific token (‘and’) from the story, the UMAP analysis shows that standard SAEs generally just cluster tokens by lexical identity. This is further corroborated by running a hierarchical clustering algorithm on the latent codes (dendrogram), finding temporally incoherent, but lexically related clusters (see Fig. 5b).
@@
Performing the experiments above on latent codes extracted from Temporal Feature Analyzers, we see that while the novel component exhibits behavior that is similar to standard SAEs’, the predictive component shows a smooth, regular curve that clusters together several tokens from the story around a similar point in space. Again running hierarchical clustering, the extracted dendrograms suggest these clusters qualitatively correspond to events occurring in the story. We emphasize this finding is very similar to the phenomenon of temporal straightening observed in neural recordings of human subjects when processing stories (xutemporal) and sequential visual data (henaff2019perceptual)! Straightening simplifies future state prediction and recent findings show that such straightening is exhibited by language models (hosseini2023large); however, Temporal Feature Analysis recovers this phenomenon in an unsupervised manner.
Quantifying Similarity to Slow vs. Fast Moving Signals. To further quantify the straightening claim, we compute the Fourier transform of the model activations and divide the frequency spectrum into two halves at a critical frequency $f_{c}$ such that the energy (i.e., sum of squared value of phase information) in the frequencies below $f_{c}$ equals that of the remaining ones. We call the first split “slow part” of a sequence, and latter the “fast part”. We then compute the correlation matrix defined by the slow and fast parts, compute their spectrum, and analyze how similar these spectrums are to the ones defined using different SAEs’ codes. Results are shown in Fig. 6. We clearly see the spectrum extracted from the predictive component of Temporal Feature Analysis approximates that of the slow part of the representation, while the novel component’s spectrum is similar to that of the fast moving part. Meanwhile, spectra of SAEs only exhibits similarity to the fast part, suggesting they do not capture longer range dependencies necessary for interpreting narrative-driven texts like most language domains. To quantify this further, we also measure kernel similarities (CKA) between the kernels respectively defined by the slow moving and fast moving signal to the different SAE codes. Results clearly show Standard SAEs are primarily similar to fast moving part of the representation.
@@
The results above demonstrate Temporal Feature Analyzers’ predictive component qualitatively align with event boundaries in a story. To investigate this result more quantitatively, we use GPT-5 to create a synthetic dataset of 50 stories with well-defined event boundaries (see Fig. 7a for an example). We extract latent codes for these stories’ tokens, center them by subtracting the mean to remove any globally shared information, and compute the cosine similarity of token to token latent codes. If the latent codes reflect local event structure of a story, the cosine similarity (on average) will be high between token pairs that come from the same event and low (if not zero) between pairs sampled across events; see Fig. 7b for an example similarity map corresponding to the story shown in Fig. 7a. Results are shown in Fig. 7 (c) and corroborate our qualitative findings: we see predictive components from Temporal Feature Analyzers show substantially higher similarity of codes if tokens are sampled from within an event, while the novel component and SAEs generally show low similarity between two tokens. These results are further supported by the robustness of Temporal Feature Analysis to noise. Specifically, we see that when we add noise to the input data, which, on average, will lead to turning off of latents with small magnitudes (due to the encoding nonlinearity), the temporal structure of the data, if it is present, will be amplified. We see precisely this effect in Fig. 7d: the predictive components’ cosine similarity map under Gaussian noised input maps elicits coarser block structures with increasing noise scale; this is reminiscent of percolation or heat diffusion perspectives on graph clustering, wherein noise diffuses only within a connected component and hence community structure is elicited (von2007tutorial). Correspondingly, we see Temporal Feature Analyzers respond most gracefully to noise: variance explained reduces slower than SAEs’, which in fact drops to $\sim$0 at some critical noise scale.
@@ 6.2 Garden Path Sentences: Ambiguities Resolved via Temporal Structure
Garden-path sentences—e.g., The old man the boat—initially cue an incorrect local parse before a later token forces reanalysis. Language models have been shown to be able to correctly parse such ambiguous sentences, offering in fact a predictive account of human per-token surprisals (li2024incremental; hanna2024incremental; oh2023transformer). Interestingly, when using SAE codes to assess whether LLM representations offer a valid parse of the sentence, we find hierarchical clustering of SAE codes yields a parse that is suggestive of the misleading cue; meanwhile, Temporal Feature Analysis recovers the correct parse by separating the predictable, slow-moving component of the representation from the novel, fast-changing residual. Specifically, in Fig. 8, we see the predictive codes link long-distance heads and dependents that reflect the correct parse (e.g., man as verb), producing coherent similarity structure over the full span, whereas standard SAEs emphasize only transient, local changes and miss these cross-temporal constraints (see App. D.4 for more examples).
@@
Figure 9: In-context representations defined by temporal relations. (a) Schematic of the synthetic “cognitive map” task defined by park2025iclrincontextlearningrepresentations: a random walk over a small graph is concatenated into a string of tokens, which when fed into a Gemma-2-2B model leads to model next-token predictions that align with the graph structure. This behavior is exhibited in the model’s activations, since a mere 2D PCA shows the graph structure. (b) Using latent codes for low-dimension projection, we see only the predictive component of the Temporal SAE clearly recovers the graph’s geometry, whereas standard SAEs and the novel component only partially reflect it. (c) Counterfactual editing: projecting the final-token representation onto the PCA basis and replacing it with a target node’s code reliably shifts next-token predictions in a graph-consistent manner.
@@ 6.3 In-Context Representations defined by entirely temporal relations
We finally use the recently described phenomenon of in-context representations by park2025iclrincontextlearningrepresentations as a domain to compare SAEs and Temporal Feature Analyzers. In particular, inspired by experiments on cognitive maps (behrens2018cognitive), park2025iclrincontextlearningrepresentations propose to take a structured graph whose nodes are assigned entities a pretrained LM is likely to have seen before during training (e.g., apple, car, etc.); see Fig. 9a for an example. A random walk of length 500 tokens is performed on this graph and the sampled nodes during the walk are concatenated together into a string. When this is inputted into a pretrained language model, provided enough context, one finds (see Fig. 9a): (i) the model begins to produce next-token predictions inline with the graph structure and (ii) a 2D PCA projection of the model activations reflects this structure. That is, the graph structure, which is randomly defined and hence impossible for the model to have seen during pretraining, is an entirely temporal construct that is inferred by the model in-context.
@@ 7.1 Geometry of latent codes in an OOD dialogue domain
We analyze a multi-turn chat domain to test whether temporal feature analysis continues to reveal slow-moving structure. We emphasize the activations in this experiment are sampled from a Gemma-2-2B-IT model, although all analyzed interpretability protocols were trained on the base (non-instruction tuned) model. Thus, in a sense, the experiments in this section also gauge the ability of different protocols to generalize out-of-distribution (OOD)—a property SAEs have been argued to partially possess when transferring between base and the corresponding instruction-tuned models (saetransfer). Specifically, we use the Ultrachat dataset, where we take the first 350 tokens from 1,000 conversations (mostly single-turn within this window, with up to four turns). For each token, we compute a shared 2D UMAP embedding from (i) raw model activations and (ii) latent codes from ReLU, TopK, and BatchTopK SAEs, as well as the novel and predictive components of the Temporal SAE. To visualize population structure clearly, we display only 100 conversations at a time to avoid clutter (Fig. 10), though fitting is applied to all 1,000.
At the population level, the predictive component of Temporal Feature Analysis is the only representation that cleanly organizes the data along two interpretable axes: (1) a speaker-role separation, with user and assistant tokens forming distinct yet smoothly adjoining manifolds, and (2) a temporal gradient, with positions flowing continuously from early to late tokens along the manifold. By contrast, standard SAEs and the novel code show diffuse clouds with weaker role separability and no coherent positional gradient (Fig. 10, top row). To assess within-sequence geometry, we overlay a single conversation’s trajectory onto the population embedding (Fig. 10, bottom row). Predictive codes trace a smooth trajectory that remains locally coherent within each role segment and transitions gracefully at speaker switches. In contrast, standard SAEs and the novel code produce jagged, zig–zagging paths. We again quantify this with tortuosity $\tau$ (bullitt2003measuring) (lower is straighter): predictive codes achieve $\tau\approx 2.0$, whereas SAEs and the novel code yield $\tau\approx 25$3030, indicating substantially more local turning (Fig. 10, panel headers). This straightening mirrors the story-domain results (Sec. 6.1), now in an OOD conversational chat regime.